OpenStudy (camerondoherty):

Question of the day: Come all who dare... What is the largest rational number less than \[\sqrt{2}\] :p

OpenStudy (solomonzelman):

1

OpenStudy (camerondoherty):

Nope

OpenStudy (solomonzelman):

Ohh, okay I meant integrer, for your question, there is no

OpenStudy (solomonzelman):

It can be any very closely approaching fraction

OpenStudy (camerondoherty):

Correct, there is none cx

OpenStudy (danjs):

none

OpenStudy (camerondoherty):

You answered too quickly :c

OpenStudy (solomonzelman):

I always think of integers, have to change my mindset.

OpenStudy (camerondoherty):

Haha

OpenStudy (unanimoose):

Is that a square of nothing? o_o

OpenStudy (solomonzelman):

I usually think of \(\large\color{black}{ \sqrt{2} }\) as of a normal number. Like, \(\large\color{black}{ \sin(45 ) =\sqrt{2}/2 }\) I see it too often to take it as something abstruse....

OpenStudy (camerondoherty):

O.o http://prntscr.com/5gx00i

OpenStudy (unanimoose):

.Sam. That's amazing.